# Visual Cue Enhancement and Dual Low-Rank Adaptation for Efficient Visual Instruction Fine-Tuning

### v2

@@ 3.4 Dual Low Rank Adaptation
The proofs of Proposition 1 and its corollary are provided in Supplementary Materials A. Inspired by these findings, we aim to explore the feasibility of using a single larger-rank LoRA to achieve the same effect mixture of LoRA experts.
3.4.1 Skill Low-Rank Space
We propose the construction of a skill low-rank space to encapsulate specific knowledge for downstream tasks while ensuring domain knowledge consistency. According to Corollary 1, a single LoRA with equivalent parameters can theoretically represent any configuration of LoRA experts across various ranks without loss of expressive power. Let $\mathcal{S}$ represent the skill space of the unified LoRA, modeled by a parameter metric $S$.
In practice, however, a split LoRA expert group often outperforms a single, larger LoRA [10]. This can be attributed to variations in routing strategies and initialization across individual LoRAs. For example, LLaVA-MoLE [4] activates the top-k experts for targeted learning, while RoDE [10] employs heterogeneous LoRA experts, each initialized with distinct distributions. The complex training process, particularly the routing score, leads to divergence in the representational spaces learned by different LoRAs over time. Intuitively, even for the same input, each expert may experience unique gradients due to its intrinsic numerical differences and routing scores.
This observation suggests the need for an additional space to mediate between the instruction task and the skill space $\mathcal{S}$, allowing $S$ to be allocated with numerical differences according to instruction variance.
3.4.2 Rectified Skill Low-Rank Space
Based on the skill low-rank space, we aim to activate prior knowledge in the task space guided by specific instructions. According to Corollary 2, additional space can be used to map the skill space $\mathcal{S}$, enabling the simulation of sparsely activated experts that respond to varying instructional tasks. We use the parameter matrix $T$ to modulate $S$, which is mapped through a non-linear activation $\sigma$ to achieve the effect of rank rectification. In this paper, we use the ReLU [11] activation function due to its sparsity-inducing properties and ease of optimization. The rectified, decomposed feature space, mapped from the skill and task spaces, can then be expressed as:
@@ 3.4.2 Rectified Skill Low-Rank Space
where $D$ denotes the Dual-LoRA module. To smooth the distribution of the skill space and improve stability in subsequent computations, we normalize the skill space using layer normalization, which can be easily implemented thanks to the unified single LoRA design. With normalization applied to the skill space, Equation 9 can be revised as:
@@ 3.5 Training
Our method consists of two training stages, as shown in Figure 1. In the vision projector pretraining stage, we train the parameters of the vision projection modules, i.e., $\{\theta_{P},\theta_{\text{deform-attn}}\}$, using descriptive image-caption pairs. In the subsequent visual instruction fine-tuning stage, we enable additional training for the Dual-LoRA parameters, where $\{\theta_{P},\theta_{\text{deform-attn}},S,T,B\}$ are set to be trainable, leveraging the full range of downstream tasks.
@@
MoE-LoRA (top-2) [27] 22.9 33.8 12.7 40.2 45.56 45.8 56.9 54.4 48.0 50.1
MoE-LoRA (softmax) [27] 22.7 33.5 12.5 40.0 45.3 45.5 58.1 53.7 47.5 50.0
@@
Table 1: Performance comparison with state-of-the-art Mixture of Experts in LoRA approaches on UniFood dataset. For fair comparison, all methods use the same base pretrained MLLM, LLaVA-1.5-7B. The best performance is highlighted in bold, and the second-best is underlined.
[…]
InstructBLIP [6] Vicuna-7B EVA-CLIP ViT-G 224 36.0 60.9 58.8 - 26.2
IDEFICS [3] LLaMA-7B OpenCLIP ViT-H/14 224 48.2 - 44.5 - -
Shikra [3] Vicuna-7B CLIP ViT-L/14 224 58.8 - - - -
Qwen-VL [1] Qwen-7B OpenCLIP ViT-bigG 448 38.2 - 62.3 - -
Qwen-VL-Chat [1] Qwen-7B OpenCLIP ViT-bigG 448 60.6 65.4 - - -
LLaVA-1.5 [15] Vicuna-7B CLIP ViT-L/14 336 64.3 - 65.4 85.9 31.1
LLaVA-1.5-LoRA [15] Vicuna-7B CLIP ViT-L/14 336 66.1 - 67.9 86.4 30.2
Dual-LoRA Vicuna-7B CLIP ViT-L/14 336 65.1 64.5 67.4 86.8 31.5
Dual-LoRA+VCE Vicuna-7B CLIP ViT-L/14 336 65.6 64.7 68.1 87.2 32.1
Table 2: Performance comparison with state-of-the-art MLLMs on general benchmarks. The best performance is highlighted in bold, and the second-best is underlined.
@@ 4 Experiments
4.1 Model and Optimizer Configuration
For the base pre-trained multimodal large language model, we use the widely adopted open-source model LLaVA-1.5-7B [14]. LLaVA-1.5-7B integrates a CLIP ViT-L [22] with an image resolution of 336px and a patch size of 14, a two-layer MLP projector to map visual features into tokens, and Vicuna v1.5 [5] as the language model. In our setup, we omit any special tokens, such as image indicator tokens, to enclose visual tokens. For the VCE module configuration, we use the outputs from the 2nd-to-last, 8th-to-last, 14th-to-last, and 20th-to-last layers of CLIP ViT-L as reference feature maps (see Section 3.3 for details). In the Dual-LoRA setting, we set the LoRA dropout rate to 0.05.
Our VCE module requires only 5.52 MB of memory, compared to the 40.02 MB needed for the vision projector in LLaVA-1.5-7B. Furthermore, our Dual-LoRA task consumes approximately 1.5 MB per LoRA rank when injected solely into the query projection and value projection layers. For optimization, we use the Adam optimizer across all trainable parameters. During the pretraining stage, we set a global batch size of 256 and a learning rate of 0.001 with no weight decay. In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LLM adapter and 0.00002 for the vision feature projector, respectively.
4.2 Evaluation on Downstream Tasks
4.2.1 Experiment Setup
@@
Dataset and Downstream Tasks. We begin by evaluating our proposed method on downstream fine-tuning tasks using the UniFood dataset [10], a multi-task dataset with 100,000 samples in the food domain. The tasks in UniFood include ingredient recognition, recipe generation, and nutrition estimation. For ingredient recognition, we use Intersection over Union (IoU) and F1-score as evaluation metrics. For recipe generation, we assess performance with SacreBLEU [21] and Rouge-L [13]. For nutrition estimation, we use the percentage of mean average error (pMAE) [32] as the metric.
@@
Baselines. We select several Mixture of Experts (MoE) methods within the LoRA framework: a sparse activation method, MoE-LoRA (top-2), which activates the top-2 experts; a dense activation method, MoE-LoRA (softmax), which uses a softmax router to engage all experts; and a rectified activation method, RoDE [10], with rectified diverse expert activation. Both MoE-LoRA (top-2) and MoE-LoRA (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$. Additionally, we include a vanilla LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
@@
Implementation Details. We conduct our experiments on a setup with 8 RTX 4090 GPUs. The LLM adapters are integrated only within the linear layers of the query and value projectors in the LLM transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.
4.2.2 Performance Comparison
Table 1 presents the results of our method compared to baseline methods on the UniFood dataset. As shown, our Dual-LoRA method consistently outperforms the sparse expert activation method (Top-2), the dense expert activation method (Softmax), and the rectified expert activation method (RoDE) across all tasks. Additionally, incorporating the proposed VCE module with Dual-LoRA (i.e., Dual-LoRA+VCE) manages to further boost performance for both ingredient recognition and nutrition estimation tasks. These results demonstrate the effectiveness of our proposed Dual-LoRA and VCE components for instruction fine-tuning in downstream tasks.
Figure 4: Qualitative results. Dual-LoRA + VCE produces more accurate and knowledge-consistent answers, whereas the LLaVA-1.5-7b-LoRA generates knowledge-inconsistent responses.
Figure 5: Performance vs. trainable LLM adapter parameters for vanilla LoRA and our proposed Dual-LoRA on the UniFood dataset. The adapters are integrated into the query and value projectors of the transformer blocks. To eliminate the influence of trainable parameters beyond the LLM adapter, only the LLM adapter is set as trainable.
4.3 Evaluation on General MLLM Benchmarks
4.3.1 Experiment Setup
MLLM Benchmarks. For the general benchmark evaluation, following LLaVA [14], we use a 558K subset of the LAION-CC-SBU dataset [23] and the llava-v1.5-mix665k dataset [14] for instruction fine-tuning. Four widely used MLLM benchmarks are used for evaluation, including MMBench [17], SEED-Bench [12], and LLaVA-Bench In-the-${}^{\text{W}}$ ) [14]. The first three benchmarks assess various MLLM capabilities, such as perceptual understanding and visual reasoning, through binary yes/no questions (MME) or multiple-choice questions (MMBench and SEED-Bench). We use the image-only subset of SEED-Bench (SEEDI), a popular choice for many image-based MLLMs [2, 30].
Implementation Details. Our experiments are conducted on a setup with 4 $\times$ A100 GPUs (80GB). The selected layers for LLM adapter injection match those used in LLaVA-1.5-7B-LoRA [14], with the adapter rank $r$ set to 128 and scaling factor $\alpha$ set to 256 in Equation 11.
4.3.2 Performance Comparison
The results of our proposed method, alongside state-of-the-art MLLMs, are presented in Table 2. The Dual-LoRA + VCE model consistently ranks among the top two across all benchmark tests. Compared to other methods, this model demonstrates robust performance across multiple benchmarks, indicating excellent generalization and strong adaptability to a variety of tasks. These findings validate the effectiveness of the VCE module and Dual-LoRA, enabling the model to achieve superior results on general visual understanding tasks.
@@ 4.3.2 Performance Comparison
Rank LMM Adapter LMM Adapter Memory Size Ingredient Recognition Recipe Generation
IoU F1 SacreBLEU Rouge-L
var1 32 LoRA 32M 19.2 29.7 10.4 41.1
var2 Dual-LoRA 48M 22.7 +18.1% 33.6 +13.3% 12.9 +24.1% 41.3 +0.3%
var3 64 LoRA 64M 20.6 31.3 12.7 41.3
var4 Dual-LoRA 96M 23.0 +11.7% 33.7 +7.7% 14.4 +13.1% 41.7 +0.9%
var5 128 LoRA 128M 21.6 32.5 13.0 41.3
var6 Dual-LoRA 192M 23.6 +8.9% 34.4 +5.9% 14.8 +14.3% 42.9 +3.6%
Table 5: Performance and memory usage of the LoRA adapter under varying rank settings on the UniFood dataset. To prevent the effect of vision projector, only the LoRA adapter is set to be trainable.
4.4 Qualitative Results
Figure 4 presents a comparison between our proposed method and the vanilla LoRA fine-tuned version of LLaVA-1.5-7B on food-related downstream tasks. The vanilla LoRA demonstrates knowledge inconsistency, with the ingredient list failing to align with the generated cooking instructions. In contrast, our method produces more accurate and knowledge-consistent answers, ensuring the ingredient list aligns with the cooking instructions. These results visually underscore the effectiveness of our approach.
@@ 4.5 Ablation Study
In this section, we conduct ablation studies to verify different components of our method on UniFood dataset. The setup is the same as Section 4.2.1.
VCE and Dual-LoRA. We conducted an ablation study to analyze the impact of different projector and adapter configurations on the performance of our model in ingredient recognition and recipe generation. The results are shown in Table 3. This ablation study shows that both multi-level local visual feature enhancement (VCE) and Dual-LoRA positively impact performance. When both are enabled, the best results are achieved for ingredient recognition and recipe generation tasks, indicating that the combination of these methods is more effective than using either one alone.
@@ 4.6 Efficiency Analysis
As listed in Table 5, we present the results of the vanilla LoRA and Dual-LoRA across three configurations of total rank values (32, 64, and 128). The results indicate that Dual-LoRA consistently outperforms the vanilla LoRA counterparts across all rank settings (32, 64, and 128) with only a slightly higher number of parameters. Notably, Dual-LoRA achieves superior performance even with fewer trainable parameters than LoRA with higher ranks, such as Var2-Dual-LoRA vs. Var3-LoRA and Var5-LoRA, as well as Var4-Dual-LoRA vs. Var5-LoRA, which is also visualized in Figure 5. We also observe that Dual-LoRA shows relatively larger improvements over vanilla LoRA when the rank size is small. We attribute this to data conflicts being more obvious with fewer trainable parameters, and the incorporation of task space in Dual-LoRA helps mitigate this issue. These findings highlight the effectiveness of Dual-LoRA in addressing data conflicts when training multiple tasks in downstream applications.
@@ 5 Conclusion
In this paper, we explore efficient visual instruction fine-tuning for multi-modal large language models (MLLMs) by leveraging pre-trained models and utilizing only a few adapter parameters. Visual instruction fine-tuning of MLLMs faces two key challenges: insufficient capture of visual cues and conflicts among various instruction tasks due to the limited adapter parameters. To address these challenges, we propose a framework comprising Vision Cues Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). For Vision cue enhancement, we use multi-level feature maps to enrich the extracted vision cues. For Dual Low-Rank Adaptation, we introduce a rectified skill-space mapping technique, enabling the model to learn complex knowledge and respond specifically to instructions. Dual Low-Rank Adaptation possesses the expressive power of any ensemble of experts with a combined rank lower than its own. Evaluations on both general visual instruction benchmarks and domain-specific benchmarks demonstrate the effectiveness of our proposed method.

### v3

@@ 3.4.1 Learning Holistic Knowledge in One Skill Space
The proof of Corollary 2 is provided in Appendix A.
3.4.2 Localized Adaptation by Skill Space Rectification
According to Corollary 2, additional space can be utilized to map $\mathcal{S}$ to simulate the “local response”, which we refer to as the task space $\mathcal{T}$. We use a parameter matrix $T$ to model $\mathcal{T}$, $T$ is mapped through a non-linear activation $\sigma$ to achieve the effect of rank rectification. In this paper, we use the ReLU [17] activation function due to its sparsity-inducing properties and ease of optimization [10, 13]. The rectified space, mapped from the skill and task spaces, can then be expressed as:
@@ 3.4.2 Localized Adaptation by Skill Space Rectification
where $D$ denotes the Dual-LoRA module. To smooth the distribution of the skill space, and improve stability in subsequent computations, we normalize the skill space using layer normalization [2], which can be easily implemented thanks to the holistic skill space design. With normalization applied to the skill space, Equation 10 can be revised as:
@@
Remark: Dual-LoRA has a much simpler structure than LoRA-MoE and offers significantly higher time efficiency, an inference time effective comparison is shown in Fig. 8.
@@ 3.5 Training
We employ a two-stage fine-tuning framework, consisting of: Stage 1: Vision Projector Pre-training, where only the VCE module and vision feature projector are trainable; and Stage 2: Visual Instruction Fine-tuning, where the VCE module, the vision projector, and the injected Dual-LoRA module are all set to trainable.
@@
LoRA-MoE (top-2) [33] 22.9 33.8 12.7 40.2 45.56 45.8 56.9 54.4 48.0 50.1
LoRA-MoE (softmax) [33] 22.7 33.5 12.5 40.0 45.3 45.5 58.1 53.7 47.5 50.0
@@
Table 1: Performance comparison with state-of-the-art Mixture of Experts in LoRA approaches on UniFood dataset. All methods use the same base pretrained MLLM, LLaVA-1.5-7B. The best performance is highlighted in bold, and the second-best is underlined.
Method ScienceQA Flickr30k
Accuracy $\uparrow$ ScareBLEU $\uparrow$ Rouge-L $\uparrow$
Vanilla LoRA 70.01 27.89 66.76
LoRA-MoE (top-2) 76.3 28.15 66.58
LoRA-MoE (softmax) 77.73 28.06 67.18
RoDE 78.39 28.17 66.89
Dual-LoRA 79.17 28.25 67.35
Dual-LoRA+VCE 80.01 28.71 67.86
Table 2: Accuracy of single-choice questions on ScienceQA and Image captioning performance on Flickr30k. The best performance is highlighted in bold, and the second-best is underlined.
@@ 4 Experiments
4.1 Experimental setup
@@
Datasets and evaluation metrics. We begin by evaluating our proposed method on downstream tasks using the following datasets and metrics: 1) UniFood [13] is a food-related dataset that includes: ingredient recognition, recipe generation, and nutrition estimation tasks. For ingredient recognition, we use Intersection over Union (IoU) and F1-score as evaluation metrics. For recipe generation, we assess performance using SacreBLEU [27] and Rouge-L [19]. For nutrition estimation, we use the percentage of mean average error (pMAE) [36] as the evaluation metric. 2) Flickr30k [26] is a large image description dataset, with each sample having 5 descriptions. All 5 descriptions are used for training to simulate potential data conflicts, SacreBLEU and Rouge-L metrics are used for evaluation. 3) ScienceQA [29] is a large-scale multimodal science question-answering benchmark covering diverse subjects. We evaluate it using single- choice questions, with average accuracy as the metric. All datasets are trained for one epoch on the training set and evaluated on the test set.
@@
Baseline methods. We select several Mixtures of LoRA Experts (LoRA-MoE) methods within the LoRA framework: a sparse activation method, LoRA-MoE (top-2) [33], which activates the top-2 experts; a dense activation method, LoRA-MoE (softmax) [33], which uses a softmax router to engage all experts; and a rectified activation method, RoDE [13], with rectified diverse expert activation. Both LoRA-MoE (top-2) and LoRA-MoE (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$. Additionally, we include a standard LoRA baseline with a rank of 64, and our Dual-LoRA is also set to rank 64.
@@
Implementation Details. We use the widely adopted open-source model LLaVA-1.5-7B [20] for the base MLLM. We conduct our experiments on a setup with 8 RTX 4090 GPUs. The LLM adapters are integrated only within the linear layers of the query and value projectors in the transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary. The detailed base model and optimization configuration can be found in Appendix B.
Hyperparameters setting. For the VCE module configuration, we select the 2nd, 8th, 14th, and 20th to the last layers of CLIP ViT-L as reference feature maps, with the 2nd layer serving as the anchor feature (see Section 3.3 for details). We set the scaling parameter $\gamma$ to 1.0 for all tasks. The VCE module requires only 5.52 MB of memory. In the Dual-LoRA setting, we set $\alpha$ to 2× the rank and the LoRA dropout rate to 0.05.
@@
4.2 Experiment Results
Performance on Multi-task Tuning. Considering the data conflicts induced by inconsistencies in optimization goals during multi-task fine-tuning, we first evaluate our approach on the multi-task datasets UniFood and ScienceQA. As shown in Table 1 for the UniFood dataset, our Dual-LoRA method consistently outperforms both vanilla LoRA and LoRA-MoE across all tasks. Similarly, for ScienceQA, as illustrated in Fig 2, our method surpasses all baseline approaches. The qualitative results in Fig. 4 further highlight the effectiveness of our approach, as it consistently generates ingredient lists across different tasks, demonstrating strong alignment with holistic domain knowledge.
Performance on Diverse Annotation Tuning. Data conflicts can also arise from diverse annotations within a single task for the same image (e.g. varying perspective descriptions of the same image), potentially distorting semantic consistency. We use Flickr30k to test our method and baseline approaches in addressing annotation diversity-induced data conflicts. As shown in Table 2, our method consistently outperforms the baseline approaches on both SacreBLEU and ROUGE-L scores, further demonstrating the effectiveness of our proposed approach. We conduct a more in-depth analysis in the following sections.
Figure 4: Conversations involving two distinct tasks can lead to knowledge-inconsistent responses with LLaVA-1.5-7b-LoRA, due to conflicts arising from multiple complex tasks. In contrast, our proposed method generates more accurate and knowledge-consistent answers.
@@ 4.3 Ablation Study
Ablation on VCE and Dual-LoRA. We conducted an ablation study to evaluate the impact of two key components: the VCE module and the Dual-LoRA module on the UniFood dataset, using the ingredient recognition and recipe generation tasks. The results are presented in Table 3. This study demonstrates that both the VCE module and the Dual-LoRA module positively influence performance. When both modules are enabled, the best results are achieved for both ingredient recognition and recipe generation tasks.
@@
Figure 5: Feature Map Visualization of Enhanced Visual Cue with VCE: (a) Enhanced Visual Cue emphasize key areas in food imagery, such as textures, garnishes, and toppings. (b) The cues highlight important details: text for readability, the cheetah-prey interaction, and the woman’s face and Oscar trophy on stage.
Figure 6: Panels (a), (b), and (c) represent feature visualizations for the recipe generation task where: (a) the distributions of feature outputs for the holistic skill space, (b) the distributions for the rectified skill space, i.e., $\mathcal{S}\odot\mathcal{T}$, and (c) the entropy of the holistic skill space (blue line) and the rectified skill space (orange line). Panels (d), (e), and (f) show the corresponding results for the nutrition estimation task.
Figure 7: Performance vs. Trainable LLM Adapter Parameters for Vanilla LoRA and Our Proposed Dual-LoRA on the UniFood Dataset. The adapters are integrated into the query and value projectors of the transformer blocks. To isolate the effect of trainable parameters beyond the LLM adapters, only the adapters are set as trainable.
Figure 8: Relative Inference Time Efficiency of Our Dual-LoRA Method and Several LoRA-MoE Methods. The inference time of vanilla LoRA is used as the baseline (unit 1). The adaptor is injected into the query and value projection layers of the LLM to improve efficiency. Test on the Flickr30k test set.
[…]
InstructBLIP [7] Vicuna-7B EVA-CLIP ViT-G 224 36.0 60.9 58.8 - 26.2
Shikra [4] Vicuna-7B CLIP ViT-L/14 224 58.8 - - - -
LLaVA-1.5 [21] Vicuna-7B CLIP ViT-L/14 336 64.3 - 65.4 85.9 31.1
LLaVA-1.5-LoRA [21] Vicuna-7B CLIP ViT-L/14 336 66.1 - 67.9 86.4 30.2
Dual-LoRA Vicuna-7B CLIP ViT-L/14 336 65.1 64.5 67.4 86.8 31.5
Dual-LoRA+VCE Vicuna-7B CLIP ViT-L/14 336 65.6 64.7 68.1 87.2 32.1
Table 5: Performance comparison between baseline methods and our proposed method on general benchmarks. The best performance is highlighted in bold, and the second-best in underlined.
4.4 Qualitative Analysis
Qualitative Results. Fig. 4 presents a comparison between the vanilla LoRA method and our proposed approach. The vanilla LoRA method exhibits knowledge inconsistency, with the ingredient list failing to align with the generated cooking instructions. In contrast, our method produces more accurate and knowledge-consistent results, where the ingredient list aligns with the cooking instructions. More qualitative results can be found in Appendix D.
Qualitative Analysis on VCE. Fig. 5 shows the heatmap of the enhanced Visual Cue, i.e., $F^{\prime}$ (see Sec. 3.3 for details). The highlighted areas closely match the regions of interest in the image, demonstrating the effectiveness of our VCE module in enhancing local visual cues.
Qualitative Analysis of Dual-LoRA. Figure 6 offers a comprehensive visualization of both the skill space and it’s rectified space, highlighting several important characteristics: 1) The information entropy of the rectified skill space is lower than that of the original skill space, as shown in panels (c) and (f) of Fig.6, suggesting that the rectified skill space is more information-determined to the current task. 2) The high-energy frequency bands in the rectified skill space exhibit regional patterns, as evidenced by the varying high values in Fig.6 (b) and (c). These observations demonstrate that the rectified skill space presents a more localized activation of the knowledge, emphasizing task-specific features.
@@ 4.5 Efficiency Analysis
Memory Usage Efficiency. Fig. 7 presents the results of vanilla LoRA and Dual-LoRA across three rank configurations (32, 64, and 128). As shown, Dual-LoRA consistently outperforms the vanilla LoRA method across all rank settings and tasks. Notably, the performance gap is more pronounced with fewer parameters. This can be attributed to the increased visibility of data conflicts when fewer parameters are trainable. Detailed data on memory usage efficiency comparisons can be found in Appendix E.
Inference Time Efficiency. Fig. 8 shows the inference time efficiency of our proposed method compared to LoRA-MoE (top-2) [33], LoRA-MoE (softmax) [33], and RoDE [13], with vanilla LoRA as the baseline (unit 1). The experimental results reveal that LoRA-MoE implementations exhibit significantly prolonged inference durations, with 2-expert and 4-expert configurations requiring 1.38× and 1.59× baseline latency respectively. In contrast, our approach demonstrates exceptional efficiency improvements: the Dual-LoRA configuration achieves only a 1.12$\times$ baseline computational overhead, and when combined with VCE, the complete system maintains high efficiency with merely a 1.16$\times$ increase in baseline latency.
4.6 Disscusion on General Benchmarks
We further conduct experiments on general benchmarks for additional analysis. The results of our proposed method, alongside state-of-the-art MLLMs, are presented in Table 5, with the experimental setup for these benchmarks detailed in Appendix B. The Dual-LoRA + VCE model consistently ranks among the top two across all benchmark tests and achieves the best performance on LLAVAW, POPE and MMVet, demonstrating the strong competitiveness of our method on general tasks.
@@ 5 Conclusion
In this paper, we explore the transition from holistic to localized adaptation for efficient visual instruction fine-tuning. We introduce Dual Low-Rank Adaptation (Dual-LoRA), which enables the model to capture holistic knowledge while being locally activated for specific tasks, thereby mitigating data conflicts. Additionally, we propose Visual Cue Enhancement (VCE), which improves the vision feature projection by incorporating local information from multi-level vision feature maps. Our methods are both time- and memory-efficient, with only a 1.16$\times$ increase in time consumption compared to standard LoRA, yet significantly outperform LoRA in terms of performance with the same parameter settings. Extensive experiments on downstream tasks validate the effectiveness of our approach.
Acknowledgements This work is supported by the Science and Technology Commission of Shanghai Municipality (Grant No. 24511103100). Additionally, it is partially supported by the Ministry of Education, Singapore, under the Academic Research Fund (AcRF) Tier 2 (Proposal ID: T2EP20222-0046). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.